A wire of radius 0.7 cm carries a current of 101 A that is uniformly distributed over its cross-sectional area. Find the magnetic field B at a distance of 0.09 cm from the center of the wire.
Answer #1

use ampere's law

Ienc = Aloop/Awire * I = I*r^2/R^2

by ampere's law B.ds = uo*Ienc

B = (uo/2pi) * (r^2/R^2) * I

here r = 0.09 x 10^-2 m = 9 x 10^-4 m

R = 0.7 cm = 7 x 10^-3 m

B = 2*10^-7 * 101 * ( 9 x 10^-4 / 4.9x 10^3.71 x 106-4 -5)

B = 3.71 x 10^-4 T